Question 3: Evaluate this model with the global test at the significance level a 0.05. (6 points) Step 1: State the hypotheses H1: Step 2: Compute the global F-statistic for the model. (Round to the nearest 100) Step 3: Find F-value for the critical value. (Round to the nearest 100) Step 4: State decision rule Step 5: State a conclusion and interpret the conclusion. Table 2 presents the parameter estimates of the regression model. Conduct a test of Question 4: the regression coefficients individually at the significance level of 0.05. Conclude which variable(s) is(are) useful in explaining the variation in Calculus grade. You are required to state the reasoning for your conclusion. Hint: the p-value approach can be used. (2 points) Table 2: Parameter Estimates Variable Intercept Algebra grade 0.961 ATC math ATC science HS rank Coefficients Standard Error Statistic p-value 36.122 0.272 0.216 0.135 10.752 0.264 0.454 3.36 3.64 0.60 0.69 01.31 0.001 0.001 0.551 0.492 0.196 0.313 0.104A professor at a college analyzed the relationship between the final grade in Calculus and factors affecting its achievement with a sample of 80 students. The independent variables included in the regression model are as follows: x1: Final grade for College Algebra, x2: ACT math score, x3: ACT natural science score, xs: Percentile high school rank. The following ANOVA summarizes the regression results. Table 1: ANOVA Source of Variation df Regression Residual or Error Total Source of Squares Mean Square F 2986.150 7,346.050 10,332.200 Question 1: (1 points) a) What is the degrees of freedom for Regression? b) What is the degrees of freedom for Residual? Question 2: (1 points) What is the value of the Regression mean square (MSR)? Round the answer to the nearest 1000th MSR- What is the value of the Error mean square (MSE)? Round the answer to the nearest a) b) 1000th MSE-
Answer #1

1)

a)

DF for regression =no. of predictors = 4

b)

DF for residuals=N-no. of predictors-1 = 80-4-1=75

2)

a) MSR=SSR/DF= 2986.15 / 4 = 746.538
b) MSE=SSE/DF = 7346.05 /75 = 97.947

3)

Anova table
variation SS df MS F-stat
regression 2986.15 4 746.538 7.6218
error, 7346.05 75 97.947
total 10332.20 79

step 1      
Ho: overall regression model is not significant      
H1: overall regression model is significant      
      
step 2      
      
F-stat=MSR/MSE =    7.62  
      
step 3      
F-critical value = 2.49 (using F-table ,alpha=0.05,df=4,75)
      
step 4-      
|F-stat|>f-critical ,then reject Ho, otherwise not      

step 5          
since F-stat > F-critical , so, rejct Ho   
hence, overall regression model is significant.

4)

for variable algebra,p-value=0.001 < alpha=0.05

so, it is significantly different from zero

for variable ATC math,p-value=0.0551>alpha=0.05, so, not significnat

fir variable ATC science,p-value = 0.492>alpha=0.05,s0 not significant

for variable, HS rank, p value=.0.196>alpha=0.05,so also not significant

so, only algebra grade is significant in explaining the variation on calculus grade